please someone help me to answer this problem, I have no idea how to do this question:4m^2+7m=0
i have no idea where to begin please help asap thanks

Answer by nerdybill(7384)

4m^2+7m=0
factor out the m:
m(4m+7)=0
.
For the above to be true:
either:
m=0
or
4m+7=0
4m=-7
m = -7/4
.
therefore, solution set:
m = {-7/4, 0}